. A person spins the pointer and is awarded the amount indicated by the pointer. pointer landed on $2
1/2 is $2, 1/4 is $8, 1/4 is $5


It costs $6 to play the game. Determine:
• The expectation of a person who plays the game.

• The fair price to play the game.

Answer by stanbon(57219)

A person spins the pointer and is awarded the amount indicated by the pointer. pointer landed on $2
1/2 is $2, 1/4 is $8, 1/4 is $5

It costs $6 to play the game. Determine:
• The expectation of a person who plays the game.
---
E(gain) = (1/2)2+(1/4)8+(1/4)5-6
---
E(gain) = 1 + 2 + 5/4 - 6
E(gain) = 4.25 - 6 = -$1.75
---
The player can expect to lose $1.75 each time he plays the game.
---
• The fair price to play the game.
The fair price would be $4.25 to play the game.
==============
Cheers,
Stan H.